Anne has a 10 3/4 pound bag of flour. Each time she bakes a batch of rolls she uses 2 1/6 pounds of flour. How many pounds of flour remain after Anne bakes 3 batches of rolls?
Answer by Boreal(15235)

10 3/4-3(2(1/6)=
3(2 1/6)=6 (1/2)
10 (3/4)-6 (1/2)=4 (1/4) pounds remain.